### v1

# Open-Set 3D object detection in LiDAR data as an Out-of-Distribution problem

Abstract
3D Object Detection from LiDAR data has achieved industry-ready performance in controlled environments through advanced deep learning methods. However, these neural network models are limited by a finite set of inlier object categories. Our work redefines the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution (OOD) problem to detect outlier objects. This approach brings additional information in comparison with traditional object detection. We establish a comparative benchmark and show that two-stage OOD methods, notably autolabelling, show promising results for 3D OOD Object Detection. Our contributions include setting a rigorous evaluation protocol by examining the evaluation of hyperparameters and evaluating strategies for generating additional data to train an OOD-aware 3D object detector. This comprehensive analysis is essential for developing robust 3D object detection systems that can perform reliably in diverse and unpredictable real-world scenarios.
I INTRODUCTION
LiDAR sensors, offering superior depth estimation and high-resolution 3D point cloud data, have become essential for autonomous vehicles, environmental mapping, and infrastructure monitoring, driving advances in 3D object detection as their availability and affordability increase. Despite significant advancements, current neural network-based 3D object detection models still encounter substantial challenges when faced with ”unknown” classes, objects that were not part of the training dataset. This problem is categorized as being ”Open-set” object detection. 3D Object detection models, while highly effective in recognizing and categorizing objects within the bounds of their training data, often struggle to generalize to new, unseen categories due to distribution shifts. Consequently, in real-world applications, where the variety and unpredictability of objects can be vast, these models can misclassify unknown objects or fail to detect them altogether. This issue is further exacerbated by the fact that the annotation of diverse object classes is both labor-intensive and costly, often leading to datasets that are not fully representative of real-world complexity. Therefore, addressing the detection of unknown classes is crucial for enhancing the robustness and reliability of 3D object detection systems in autonomous vehicles and other applications reliant on LiDAR data.
Specifically, traditional 3D object detection aims at outputting, for a given 3D LiDAR scan, bounding boxes represented as follows:
where , and represent the center of the bounding box along the and coordinates, and represent the length, width and height, and represents the heading of the bounding box, and represents the predicted class amongst the annotated classes in the dataset. These bounding boxes should overlap with the ground truth bounding boxes. In Open-Set 3D Object Detection, the output is similar, but now represents the predicted class amongst the annotated classes in the dataset and an additional ”unknown” class, representing all other relevant objects in the scenes. Another similar and relevant problem is Out-Of-Distribution (OOD) Detection, where from a set of predictions, an out-of-distribution measure is added:
Works focusing on Open-Set 3D Detection suffer from a model performance degradation on known classes [openset3Dobj, Liang2023UnknownSF, towardsopen, mc_dropout]. Framing the problem as an out-of-distribution task allows us to maintain the original model performance.
In this paper, we redefine the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution (OOD) problem, setting a rigorous evaluation protocol by examining evaluation hyperparameters. Additionally, we evaluate methods for generating additional data to train an OOD-aware 3D object detector.
We summarize our contributions as follows:
Redefining the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution problem..
Setting a rigorous evaluation protocol through examining evaluation hyperparameters.
Evaluating generation methods of additional data for training an OOD-aware 3D object detector.
II Related Works
II-A 3D Object Detection
Recent advancements in 3D object detection have led to the development of several notable architectures that enhance the accuracy and efficiency of LiDAR-based perception systems. VoxelNet [Zhou_Tuzel_2018] introduced the concept of end-to-end learning directly from raw point cloud data by voxelizing the space, which significantly improved feature extraction capabilities. Building on this, PointPillars [Lang_2019_CVPR] streamlined the voxelization process by converting point clouds into pseudo-images, enabling faster and more efficient processing suitable for real-time applications. CenterPoint [Yin_2021_CVPR] further improved detection accuracy by proposing a center-based representation, simplifying object localization. Additionally, DSVT [wang2023dsvt] introduced a transformer-based approach to dynamically process sparse voxel grids, enhancing the detection system’s robustness and scalability.
II-B Open-Set and Out-Of-Distribution Detection
Several works have advanced open-set detection in 2D out-of-distribution (OOD) detection. VOS [du2022vos] introduced a technique to synthesize virtual outliers, improving a model’s capability to distinguish between known and unknown classes by exposing it to artificially generated outlier examples during training. However, the method contains many hyperparameters for defining class-conditional Gaussian distributions, and they find that fixed Gaussian noise achieves close results. ODIN [odin] enhanced confidence estimates for OOD detection using temperature scaling and input perturbation, leading to more reliable detection of OOD samples. MSP [msp] proposed using the maximum softmax probability as a baseline for OOD detection. Liu et al. [energy] leveraged neural network output energy scores to differentiate between in-distribution and OOD samples. [hendrycks17baseline] provided a robust baseline by analyzing softmax scores to detect misclassified and OOD examples. On the other hand, ODIN, MSP, and Energy thresholding are based on in-distribution data and lack any integration of the semantic information of unknown objects. Finally, [towardsopen] addressed the broader challenge of dynamically recognizing new, unseen classes, proposing methodologies to handle an evolving set of object categories in open-world scenarios.
In 3D object detection, handling OOD scenarios has also seen substantial progress. Cen et al. [openset3Dobj] explored methods for detecting and managing unknown objects within 3D environments. However, their evaluations are based on UDI, a private dataset with only 600 unknown test objects, and KITTI, which we will show is insufficient. UFO [UFO] introduced novel approaches for identifying unknown objects in point clouds, focusing on accurately segmenting foreground objects, but they also focus on the limited KITTI dataset. Huang et al. [Huang2022OutofDistributionDF] explored different instance injection methods to test OOD baselines. The variance in their results underlines the difficulty in identifying the OOD objects they insert from different sources, which have different representations from the KITTI dataset in which they are injected. Lastly, Kösel et al. [revisiting] offered an evaluation of existing OOD techniques in the context of LiDAR data, identifying key challenges and proposing enhancements. However, their work lacks specifications on their evaluation protocol and evaluated objects, such as the recall/hit percentage of unknowns, and only test resizing unknown generation on nuScenes, which has certain limitations as we will see in Section III. We aim to improve the evaluation methodology behind their work, set a rigorous benchmark of unknown generation, and link it to the global aim of open-set detection. We find that using a focal loss instead of a binary cross-entropy loss yields better results. Furthermore, their resizing unknown generation method produces lesser results than top-K labelling, suggesting that using pseudo-labels from real OOD data can lead to better performance than reshaping known objects.
III Revisiting Open-Set Detection
III-A Viewing open-set 3D object detection as an OOD problem
Closed-set object detectors are optimized and trained to detect a set of ”core” closed classes that are essential for autonomous vehicle tasks.
Previous works [openset3Dobj, Liang2023UnknownSF, towardsopen, mc_dropout] have focused on detecting unknown objects as a new, separate class at the expense of the base detector’s performance on known classes. We can view the Open Set detection as an out-of-distribution (OOD) prediction challenge, where our focus is on evaluating a base detector’s predictions and determining which of these correspond to genuine unknown objects in the scene. This allows the detection output to be enhanced with additional information, the known class, which is semantically closest to the unknown object, along with an anomaly score. Other modules could use this information in an autonomous vehicle pipeline, which is not the case for Open-set object detection, in which ”unknown” is considered a separate class entirely. We illustrate this difference in Figure 1.
This suggests we could view the open-set problem through an out-of-distribution approach, enhancing the base detector’s ability to handle open-set scenarios without compromising its performance on the core closed classes.
This perspective sets the stage for discussing the strategies essential for evaluating out-of-distribution detection models.
III-B Datasets for OOD Evaluation
III-B1 The limits of widely used datasets
Currently, when addressing the out-of-distribution (OOD) detection problem in 3D object detection, the limitations of existing datasets such as KITTI [Geiger2012CVPR], ONCE [once_ds], and Waymo [Sun_2020_CVPR] become particularly pronounced. Indeed, these datasets contain a small number of classes, respectively 5, 5, and 3. Methods such as [openset3Dobj], which evaluate on the Van and Truck subset of KITTI, tackle a specific part of the issue, these classes being both vehicles and lacking the high diversity which can arise in real-world scenarios.
nuScenes [caesar2020nuscenes], while one of the most comprehensive datasets, presents specific challenges critical in the context of OOD detection. Firstly, the predefined dataset splits and evaluation metrics in nuScenes might not be sufficient for evaluating OOD performance. For instance, only one out of four of the dataset’s scans in its evaluation set contains unknown objects as defined by [revisiting], thus limiting the evaluation interpretation. nuScenes also suffers from its point cloud sparsity due to its 32-beam LiDAR sensor, which can only be solved by aggregating scans and creating a trailing point cloud for moving objects, hindering detection. These findings underscore the need to use a more diverse and dense dataset to effectively tackle the OOD problem in 3D object detection.
III-B2 Argoverse 2
In order to establish a robust baseline for out-of-distribution (OOD) detection, we turn to the Argoverse 2 dataset [wilson2021argoverse], which offers several distinct advantages over other datasets such as nuScenes. Argoverse 2 is noted for its high class diversity, encompassing a wide range of object types and scenarios, which is crucial for training models that can generalize well to new, unseen objects. This dataset also provides better LiDAR representation, with higher density point clouds than nuScenes, due to its use of two 32-beam LiDAR sensors, without the ”point cloud trail” phenomenon. The dataset contains scans in the training set and in its validation set.
For Argoverse 2, we choose the rarest 5% of classes as our open classes as in [revisiting], denoted as set . This approach ensures that the chosen open classes are genuinely underrepresented in the training data, providing a challenging test for the OOD detection capabilities of our models. This selection method helps simulate real-world scenarios where we assume that the most common objects will have been labeled beforehand while certain rare objects are infrequently encountered, making it a valuable benchmark for assessing the robustness of 3D object detection systems. The open classes , representing instances in the validation set, are:
| Articulated bus, Stroller, Motorcycle, Mobile pedestrian crossing | |||
The closed ”known” classes, denoted as set , include more common objects, presumed to have been labeled beforehand. These classes, representing instances in the validation set, are:
| Stop sign, Sign, Bus, Truck, Bicycle, Bicyclist, Wheeled device, | |||
III-B3 Metrics and Matching
In our evaluation, we use metrics commonly employed in OOD detection, including AUROC, FPR-95, AUPR-E (with open classes as positives), and AUPR-S (with known classes as positives), to assess the model’s ability to distinguish between known and unknown objects. We also describe our matching algorithm, detailed in Algorithm 1, which is crucial for computing True Positives (TP), False Positives (FP), True Negatives (TN), and False Negatives (FN). This algorithm, which includes essential hyperparameters, plays a significant role in determining the final evaluation metrics, mainly by ignoring any predicted box whose center is not within a fixed range of any non-assigned ground truth.
III-C Single and Two-Stage Methods for OOD
Methods to compute ”unknown” object scores can be generally separated into two categories: single-stage methods, which directly use the outputs of a detector model and apply some thresholding or entropy computation, and two-stage methods, which add additional trainable parameters on top of the base detector. This latter approach necessitates a retraining phase. We illustrate these two families of methods in Figure 2.
III-C1 Single-Stage Methods
In our evaluation of OOD detection for 3D object detection, we selected several single-stage methods to ensure a comprehensive assessment, i.e. methods that directly use the output logits of the detection model. The chosen methods include ODIN [odin], MaxLogit [msp], Energy Thresholding [energy], MC Dropout inspired by [mc_dropout], and the Default Score.
III-C2 Two-Stage Methods
Like [revisiting], we integrated an additional Multi-Layer Perceptron (MLP) module after the base detector. This module is specifically trained to output a binary classification confidence score or an ”unknown” confidence score: 0 for known objects and 1 for unknown objects. This two-stage solution employs the MLP to process initial detections, decomposed into embeddings, box predictions, and prediction logits and outputs an ”out-of-distribution” score . During inference, if is larger than a given threshold , the object is considered an OOD object.
III-D Analysis of the Evaluation Protocol for 3D OOD
We use the ground truth data to train the MLP with the actual annotated ’unknown’ classes, which we will refer to as the ’Oracle MLP.’ We demonstrate that many external factors could impact the performance and resulting metrics, which are not disclosed in [revisiting]. We identify several key parameters that significantly influence the effectiveness of the Oracle MLP when applied to GT data, allowing us to examine the influence of these parameters independently of the actual performance of the MLP.
In the first stage, we train a CenterPoint [Yin_2021_CVPR] object detector on the Argoverse 2 dataset [wilson2021argoverse] with classical training regimes [revisiting, Yin_2021_CVPR]. This model obtained a mAP score of 35.2 on the known classes. For the two-stage family of methods, we train an MLP after freezing the initial detector Centerpoint. We use a 3-layer MLP where we divide the dimensionality by two between every layer, followed by a sigmoid activation function, on top of our frozen pretrained CenterPoint model. We train it for 5 epochs, with a polynomial learning rate using a power of , with an initial learning rate of . For ease of reading, we include relevant experiments from the best single-stage model ODIN [odin] (see Section IV-A) and the two-stage Oracle MLP in the following.
By systematically analyzing the evaluation hyperparameters, we aim to provide a more comprehensive understanding of how different evaluation factors influence the final performance of unknown OOD detection. These hyperparameters include:
III-D1 Evaluation set
Using the ”Open” Scans subset, which includes only scans with unknown objects, significantly improved the AUPR-E metric from 12.6 to 20.0 for the 2-stage Oracle MLP. Due to the known/unknown class imbalance, using all scans adds many more potential false positives without any unknown ground truth, hindering all methods across the board. Consequently, our final evaluation used the ”open” subset—only the evaluation scans containing at least one unknown object—to better assess performance.
III-D2 Distance and Threshold Settings for OOD Detection
The choices of distance metric and detection module threshold play a critical role in OOD Detection.
Adjusting the distance threshold for matching predicted boxes to ground truth objects influences the inclusion of predictions in metric calculations. Specifically, when using a larger (e.g., 2.0m instead of 0.5m), more predictions, mainly those correctly identifying unknown classes, are included. This increase in detection hits—predicted boxes within the range of ground truth—allows for a more comprehensive evaluation of the model’s ability to detect unknown objects. However, this also introduces a risk of misassigning predictions to known objects, potentially lowering AUROC and FPR-95 while improving AUPR-E, as this metric directly reflects the precision in detecting unknowns.
A lower threshold in the detection module (e.g., 0.3) allows more predicted objects to pass through, reducing misses, especially for known objects. However, this comes at the cost of an increased likelihood of false positives among the unknown detections. A trade-off exists: while a lower threshold increases the detection rate of unknown objects, it also results in more potential false positives. Therefore, a threshold of 0.3, paired with a 2.0m distance for box matching, strikes a balance between minimizing misses for known objects and enhancing the detection of unknown objects, as evidenced by the metrics in Table I.
| Hits (O)% | Hits (C)% | Method | AUROC | FPR-95 | AUPR-E | AUPR-S | ||
| 0.05 | 0.5m | 22.8 | 49.3 | ODIN | 88.3 | 62.4 | 10.1 | 97.3 |
| Oracle MLP | 91.0 | 43.7 | 22.2 | 99.8 | ||||
| 2.0m | 38.9 | 55.4 | ODIN | 82.1 | 78.6 | 9.1 | 97.4 | |
| Oracle MLP | 84.9 | 62.8 | 17.5 | 99.5 | ||||
| 0.3 | 0.5m | 14.4 | 45.3 | ODIN | 89.9 | 47.8 | 9.8 | 96.9 |
| Oracle MLP | 92.3 | 37.4 | 23.2 | 99.9 | ||||
| 2.0m | 26.2 | 50.6 | ODIN | 85.9 | 61.6 | 9.8 | 97.1 | |
| Oracle MLP | 88.6 | 51.1 | 20.0 | 99.7 | ||||
| 0.5 | 0.5m | 7.6 | 39.0 | ODIN | 91.4 | 38.0 | 17.0 | 96.4 |
| Oracle MLP | 93.6 | 32.6 | 24.8 | 99.9 | ||||
| 2.0m | 11.7 | 42.4 | ODIN | 87.4 | 48.9 | 14.2 | 96.5 | |
| Oracle MLP | 91.4 | 38.7 | 22.1 | 99.9 |
III-D3 Sorting method
In our experiments, we observed a 2.0 increase in the AUROC metric when sorting predictions by the base detector’s confidence score rather than by the ”unknown” score confidence. This suggests that using the original detector’s score is more reliable for prioritizing predictions, especially given the high prediction scores for unknown objects and the non-overlap hypothesis in 3D space.
From Table I, we notice widely different results from the traditional OOD metrics for any fixed method, depending on the evaluation hyperparameters. Notably, they modify the performance differences between methods. For our following evaluations, we decided to strike a balance between a high recall of unknowns, which includes noisy predictions and complex examples, and a lower recall, which only isolates well-defined examples. We suggest that future works take into account the recall of the base detector and specify these important evaluation hyperparameters.
IV Comparison of single, two-stage methods and data generation for OOD training
IV-A Limits of single-stage methods
We compared multiple single-stage methods for ODD detection on the Argoverse 2 dataset with our evaluation protocol of Section III. Table II shows the results. All single-stage models fail to reach the performance of the 2-stage oracle MLP. While the ODIN method achieves the highest AUROC (85.9 on Argoverse 2), it still exhibits a high False Positive Rate at 95% recall (61.6%). Energy-based methods, Monte-Carlo Dropout, MaxLogit, and Default Score thresholding yield relatively weaker results. These consistently mediocre performances across all methods indicate that single-stage approaches are lacking compared to a model trained directly on unknown object representations. Therefore, we experimented with two-stage methods to try and match this oracle.
| Method | AUROC | FPR-95 | AUPR-E | AUPR-S |
| Default Score | 76.6 | 86.5 | 4.3 | 99.4 |
| Energy [energy] | 65.8 | 94.6 | 2.7 | 99.0 |
| Monte Carlo Dropout [mc_dropout] | 75.4 | 69.7 | 7.0 | 98.9 |
| MaxLogit [msp] | 76.6 | 86.5 | 4.3 | 98.6 |
| ODIN [odin] | 85.9 | 61.6 | 9.8 | 97.1 |
| 2-Stage Oracle MLP | 86.8 | 51.1 | 20.0 | 99.7 |
IV-B Two-Stage Methods and Input Choices
The two-stage methods use an MLP for OOD prediction. We first compare different inputs of the MLP using GT data (named ’Oracle MLP’ as in Section III-D):
high-dimensional feature map (High-Dim): we can use higher-dimension (512) feature map representation from the base detector instead of the final lower-dimension (192) feature maps
feature interpolation (Interp.): Instead of extracting the feature representation of the base detector’s boxes at the discrete grid value of the feature map, we instead interpolate these features based on the predicted center
using a focal loss instead of a binary cross-entropy loss (F. Loss)
using a max-pooling (Pool) of the feature map before extracting the feature vector at the corresponding feature map location of the predicted object center
| High-Dim | Interp. | F. Loss | Pool | AUROC | FPR-95 | AUPR-E | AUPR-S |
| 85.2 | 52.6 | 17.6 | 70.4 | ||||
| ✓ | 77.3 | 66.9 | 10.2 | 79.1 | |||
| ✓ | 85.4 | 51.6 | 17.8 | 70.6 | |||
| ✓ | 86.8 | 50.6 | 15.7 | 99.7 | |||
| ✓ | 88.7 | 44.5 | 26.5 | 71.8 | |||
| ✓∗ | ✓∗ | 88.6 | 43.2 | 27.1 | 70.5 | ||
| ✓ | ✓ | ✓ | ✓ | 88.6 | 51.1 | 20.0 | 99.7 |
We show our experiments using the ground truth unknown objects for training in Table III. The results show that adding a 3x3 pooling layer and a focal loss was beneficial, compared to the work [revisiting], which only uses a binary cross-entropy loss, no pooling, high-dimensional feature maps, and feature interpolation. Despite a higher FPR-95, this is the only method that combines a higher precision on known objects and near-perfect precision on closed objects, which was outlined as a primary objective when seeing Open-Set as an OOD problem. We highlight this in the table, where using the input processing of [revisiting] leads to a known object precision of 70.5, and ours a 99.7 precision while maintaining an acceptable AUPR-E. The pooling smooths out the center interpolation errors from the detection model. For instance, in the Argoverse 2 dataset, a maxpooling filter enabled a feature aggregation operation with a radius of 2.1 meters. It substantially improved OOD performance by ensuring a more robust aggregation of surrounding point cloud data. While we trained our MLP models using the ground truth unknown objects in the previous section, we cannot access this information when training in real-world scenarios. Therefore, we need to generate pseudo-labels to train our model.
IV-C Methods for Generating Additional Unknown data for training the MLP
We then trained our MLP to detect OOD objects, using different generation methods of unknown objects, which we use as our ground truth.
IV-C1 Different Representations of Out-of-Distribution Objects
We compare different representations of out-of-distribution objects for training, which we create through several means:
Gaussian noise [du2022vos]: we perturb the extracted embedding of ID objects with Gaussian noise, adding a normal distribution to the extracted embeddings.
External data: we incorporate ModelNet [modelnet], allowing us to introduce a diverse set of synthetic objects into our training regime, which we consider OOD. We insert between 15 and 25 objects, sampled randomly from the {’bench’, ’bookshelf’, ’chair’, ’cone’, ’desk’, ’flower pot’, ’lamp’, ’mantel’, ’night stand’, ’plant’, ’vase’, ’wardrobe’} class subset of ModelNet, and normalize their size to 1. We then sample 200 points at the surface of each object, followed by a grid sampling where we keep between 1 and 3 points with a grid size which we also randomize between 5 and 10, apply an intensity value to the points based on the intensity statistics of Argoverse 2, and finally resize the objects with a global scaling factor between 1 and 4.
PointMixup [pointmixup]: we use this method to mix point clouds of different objects through point matching and interpolation to create new, hybrid representations. We expect objects created in such a fashion to contain local features present in both ID and OOD objects while creating novel appearances. Since [pointmixup] was designed for data augmentation, we chose different interpolation hyperparameters after testing variations. Each object with at least 5 points has a 20% probability of being replaced by a new version of that object generated as an interpolation between it and another object with at least 5 points, with interpolation factors randomly chosen in the range.
Resizing: create out-of-distribution objects by resizing in-distribution (ID) known objects to modify their scales along all three axes, similarly to [revisiting].
Top-K autolabelling approach [towardsopen]: the top-K most confident predictions from the base detector, which don’t overlap with ground truth ID objects, are used as OOD objects. In this approach, we keep the scans from the datasets containing real OOD distribution, as we expect these to become autolabelled. Since other methods do not rely on real OOD data for training the MLP, we remove those scans for other methods.
We illustrate these different generation methods in Figure 3 and compare these various representations of OOD objects and their impact on the training and performance of our additional module on real validation OOD data.
IV-C2 Results
The results of our evaluation on the Argoverse 2 dataset, presented in Table IV, highlight the discrepancies in the MLP performance based on different generation methods of unknowns.
The Resizing method obtained an AUROC of 77.2, with a relatively low AUPR-E of 6.3. Compared to the original paper, these low results highlight how this simple approach might not be suitable for the more diverse Argoverse 2. The Gaussian Noise method performed poorly across all metrics, with an AUROC of 66.5 and a very high FPR-95 of 84.5, making it only slightly better than random chance.
| Method | AUROC | FPR-95 | AUPR-E | AUPR-S |
| Gaussian Noise [du2022vos] | 66.5 | 84.5 | 3.7 | 98.9 |
| ModelNet Injection† | 70.6 | 78.7 | 5.0 | 71.1 |
| PointMixUp†[pointmixup] | 72.1 | 72.4 | 7.4 | 99.1 |
| Resizing [revisiting] | 77.2 | 75.3 | 6.3 | 99.4 |
| Top-K Labelling (K=5) [towardsopen] | 81.9 | 68.2 | 8.8 | 99.6 |
| 2-Stage Oracle MLP | 86.8 | 51.1 | 20.0 | 99.7 |
ModelNet Injection, which incorporates synthetic objects from the ModelNet dataset as OOD data, showed better performance with an AUROC of 70.6 and an AUPR-E of 5.0. We believe the difficulties in OOD detection for this method stem from the synthetic nature of the dataset, which suggests that while other works suggest using insertions of instances as unknowns [UFO], one has to be careful about the quality of these instances and the domain differences between these and the real unknowns. PointMixUp struggled with a high FPR-95 of 77.6 and a low AUPR-E of 4.6, indicating challenges in detecting unknown objects despite the novel hybrid representations. The Top-K Labelling (K=5) method, which labels the top K most confident predictions not overlapping with ground truth ID objects as OOD, emerged as a notable performer. It achieved an AUROC of 81.9 and the highest AUPR-S of 99.6, demonstrating exceptional identification of known objects. Despite its higher FPR-95 of 68.2, Top-K Labelling showed promise in balancing the detection of known and unknown objects by leveraging real-world OOD distributions in the dataset. We experimented with different K values (K= ) and found that K=5, close to the average number of open class instances per frame, performed best across all metrics. Overall, the remaining gap between the two-stage oracle MLP and this highest-performing unknown generation process suggests that future works should focus on improving this pseudo-labelling process.
V CONCLUSIONS
Our work redefines the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution problem. Through this lens, OOD 3D object detection adds relevant information with regard to traditional object detection. In the process, we established a rigorous evaluation protocol and identified key evaluation parameters that we suggest highlighting in future works. Additionally, we show that two-stage methods show promise for OOD 3D Object Detection and underline the potential of autolabelling methods.

### v2

@@
-Open-Set 3D object detection in LiDAR data as an Out-of-Distribution problem
+HD-OOD3D: Supervised and Unsupervised Out-of-Distribution object detection in LiDAR data
 Abstract
-3D Object Detection from LiDAR data has achieved industry-ready performance in controlled environments through advanced deep learning methods.
-However, these neural network models are limited by a finite set of inlier object categories.
-Our work redefines the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution (OOD) problem to detect outlier objects.
-This approach brings additional information in comparison with traditional object detection.
-We establish a comparative benchmark and show that two-stage OOD methods, notably autolabelling, show promising results for 3D OOD Object Detection.
-Our contributions include setting a rigorous evaluation protocol by examining the evaluation of hyperparameters and evaluating strategies for generating additional data to train an OOD-aware 3D object detector.
-This comprehensive analysis is essential for developing robust 3D object detection systems that can perform reliably in diverse and unpredictable real-world scenarios.
+Autonomous systems rely on accurate 3D object detection from LiDAR data, yet most detectors are limited to a predefined set of known classes, making them vulnerable to unexpected out-of-distribution (OOD) objects.
+In this work, we present HD-OOD3D, a novel two-stage method for detecting unknown objects.
+We demonstrate the superiority of two-stage approaches over single-stage methods, achieving more robust detection of unknown objects while addressing key challenges in the evaluation protocol.
+Furthermore, we conduct an in-depth analysis of the standard evaluation protocol for OOD detection, revealing the critical impact of hyperparameter choices.
+To address the challenge of scaling the learning of unknown objects, we explore unsupervised training strategies to generate pseudo-labels for unknowns.
+Among the different approaches evaluated, our experiments show that top-5 auto-labelling offers more promising performance compared to simple resizing techniques.
 I INTRODUCTION
@@
 Despite significant advancements, current neural network-based 3D object detection models still encounter substantial challenges when faced with ”unknown” classes, objects that were not part of the training dataset.
-This problem is categorized as being ”Open-set” object detection.
-3D Object detection models, while highly effective in recognizing and categorizing objects within the bounds of their training data, often struggle to generalize to new, unseen categories due to distribution shifts.
-Consequently, in real-world applications, where the variety and unpredictability of objects can be vast, these models can misclassify unknown objects or fail to detect them altogether.
+In real-world scenarios, encountering objects that lie outside the training distribution—referred to as out-of-distribution (OOD) objects—is inevitable and can pose significant safety risks.
 This issue is further exacerbated by the fact that the annotation of diverse object classes is both labor-intensive and costly, often leading to datasets that are not fully representative of real-world complexity.
-Therefore, addressing the detection of unknown classes is crucial for enhancing the robustness and reliability of 3D object detection systems in autonomous vehicles and other applications reliant on LiDAR data.
 Specifically, traditional 3D object detection aims at outputting, for a given 3D LiDAR scan, bounding boxes represented as follows:
@@
 Another similar and relevant problem is Out-Of-Distribution (OOD) Detection, where from a set of predictions, an out-of-distribution measure is added:
-Works focusing on Open-Set 3D Detection suffer from a model performance degradation on known classes [openset3Dobj, Liang2023UnknownSF, towardsopen, mc_dropout].
-Framing the problem as an out-of-distribution task allows us to maintain the original model performance.
-In this paper, we redefine the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution (OOD) problem, setting a rigorous evaluation protocol by examining evaluation hyperparameters.
-Additionally, we evaluate methods for generating additional data to train an OOD-aware 3D object detector.
+Works focusing on Open-Set 3D Detection suffer from a model performance degradation on known classes [1, 2, 3, 4], which is why we frame the problem as an out-of-distribution task.
+The latter also preserves the original semantic information of the predicted object which is necessary for downstream tasks.
+A central component of our work is the development of a rigorous evaluation protocol.
+We show that key hyperparameters, such as distance thresholds and detection score cut-offs, have a substantial impact on the performance of OOD detection.
+By systematically analyzing these factors, we provide a more comprehensive and reproducible framework for evaluating OOD-aware detectors.
+Furthermore, we investigate a variety of strategies for generating additional training data - from synthetic data injections to autolabelling techniques - to improve the model’s ability to recognize and handle OOD objects.
 We summarize our contributions as follows:
-Redefining the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution problem..
-Setting a rigorous evaluation protocol through examining evaluation hyperparameters.
+Designing a new method for out-of-distribution 3D object detection.
+Setting a rigorous evaluation protocol by examining evaluation hyperparameters.
 Evaluating generation methods of additional data for training an OOD-aware 3D object detector.
@@
 Recent advancements in 3D object detection have led to the development of several notable architectures that enhance the accuracy and efficiency of LiDAR-based perception systems.
-VoxelNet [Zhou_Tuzel_2018] introduced the concept of end-to-end learning directly from raw point cloud data by voxelizing the space, which significantly improved feature extraction capabilities.
-Building on this, PointPillars [Lang_2019_CVPR] streamlined the voxelization process by converting point clouds into pseudo-images, enabling faster and more efficient processing suitable for real-time applications.
-CenterPoint [Yin_2021_CVPR] further improved detection accuracy by proposing a center-based representation, simplifying object localization.
-Additionally, DSVT [wang2023dsvt] introduced a transformer-based approach to dynamically process sparse voxel grids, enhancing the detection system’s robustness and scalability.
+VoxelNet [5] introduced the concept of end-to-end learning directly from raw point cloud data by voxelizing the space, which significantly improved feature extraction capabilities.
+Building on this, PointPillars [6] streamlined the voxelization process by converting point clouds into pseudo-images, enabling faster and more efficient processing suitable for real-time applications.
+CenterPoint [7] further improved detection accuracy by proposing a center-based representation, simplifying object localization.
+Additionally, DSVT [8] introduced a transformer-based approach to dynamically process sparse voxel grids, enhancing the detection system’s robustness and scalability.
 II-B Open-Set and Out-Of-Distribution Detection
 Several works have advanced open-set detection in 2D out-of-distribution (OOD) detection.
-VOS [du2022vos] introduced a technique to synthesize virtual outliers, improving a model’s capability to distinguish between known and unknown classes by exposing it to artificially generated outlier examples during training.
+VOS [9] introduced a technique to synthesize virtual outliers, improving a model’s capability to distinguish between known and unknown classes by exposing it to artificially generated outlier examples during training.
 However, the method contains many hyperparameters for defining class-conditional Gaussian distributions, and they find that fixed Gaussian noise achieves close results.
-ODIN [odin] enhanced confidence estimates for OOD detection using temperature scaling and input perturbation, leading to more reliable detection of OOD samples.
-MSP [msp] proposed using the maximum softmax probability as a baseline for OOD detection.
+ODIN [10] enhanced confidence estimates for OOD detection using temperature scaling and input perturbation, leading to more reliable detection of OOD samples.
+MSP [11] proposed using the maximum softmax probability as a baseline for OOD detection.
 Liu et al.
-[energy] leveraged neural network output energy scores to differentiate between in-distribution and OOD samples.
-[hendrycks17baseline] provided a robust baseline by analyzing softmax scores to detect misclassified and OOD examples.
+[12] leveraged neural network output energy scores to differentiate between in-distribution and OOD samples.
+[13] provided a robust baseline by analyzing softmax scores to detect misclassified and OOD examples.
 On the other hand, ODIN, MSP, and Energy thresholding are based on in-distribution data and lack any integration of the semantic information of unknown objects.
-Finally, [towardsopen] addressed the broader challenge of dynamically recognizing new, unseen classes, proposing methodologies to handle an evolving set of object categories in open-world scenarios.
+Finally, [3] addressed the broader challenge of dynamically recognizing new, unseen classes, proposing methodologies to handle an evolving set of object categories in open-world scenarios.
 In 3D object detection, handling OOD scenarios has also seen substantial progress.
 Cen et al.
-[openset3Dobj] explored methods for detecting and managing unknown objects within 3D environments.
+[1] explored methods for detecting and managing unknown objects within 3D environments.
 However, their evaluations are based on UDI, a private dataset with only 600 unknown test objects, and KITTI, which we will show is insufficient.
-UFO [UFO] introduced novel approaches for identifying unknown objects in point clouds, focusing on accurately segmenting foreground objects, but they also focus on the limited KITTI dataset.
+UFO [14] introduced novel approaches for identifying unknown objects in point clouds, focusing on accurately segmenting foreground objects, but they also focus on the limited KITTI dataset.
 Huang et al.
-[Huang2022OutofDistributionDF] explored different instance injection methods to test OOD baselines.
+[15] explored different instance injection methods to test OOD baselines.
 The variance in their results underlines the difficulty in identifying the OOD objects they insert from different sources, which have different representations from the KITTI dataset in which they are injected.
 Lastly, Kösel et al.
-[revisiting] offered an evaluation of existing OOD techniques in the context of LiDAR data, identifying key challenges and proposing enhancements.
-However, their work lacks specifications on their evaluation protocol and evaluated objects, such as the recall/hit percentage of unknowns, and only test resizing unknown generation on nuScenes, which has certain limitations as we will see in Section III.
-We aim to improve the evaluation methodology behind their work, set a rigorous benchmark of unknown generation, and link it to the global aim of open-set detection.
+[16] offered an evaluation of existing OOD techniques in the context of LiDAR data, identifying key challenges and proposing enhancements.
+However, their work lacks specifications on their evaluation protocol and evaluated objects, such as the recall/hit percentage of unknowns, and only test resizing unknown generation on nuScenes, which has certain limitations as we will see in Section IV.
+We aim to improve the evaluation methodology behind their work as well as the performance through a rigorous benchmark of unknown generation.
 We find that using a focal loss instead of a binary cross-entropy loss yields better results.
 Furthermore, their resizing unknown generation method produces lesser results than top-K labelling, suggesting that using pseudo-labels from real OOD data can lead to better performance than reshaping known objects.
-III Revisiting Open-Set Detection
-III-A Viewing open-set 3D object detection as an OOD problem
-Closed-set object detectors are optimized and trained to detect a set of ”core” closed classes that are essential for autonomous vehicle tasks.
-Previous works [openset3Dobj, Liang2023UnknownSF, towardsopen, mc_dropout] have focused on detecting unknown objects as a new, separate class at the expense of the base detector’s performance on known classes.
-We can view the Open Set detection as an out-of-distribution (OOD) prediction challenge, where our focus is on evaluating a base detector’s predictions and determining which of these correspond to genuine unknown objects in the scene.
-This allows the detection output to be enhanced with additional information, the known class, which is semantically closest to the unknown object, along with an anomaly score.
-Other modules could use this information in an autonomous vehicle pipeline, which is not the case for Open-set object detection, in which ”unknown” is considered a separate class entirely.
-We illustrate this difference in Figure 1.
-This suggests we could view the open-set problem through an out-of-distribution approach, enhancing the base detector’s ability to handle open-set scenarios without compromising its performance on the core closed classes.
-This perspective sets the stage for discussing the strategies essential for evaluating out-of-distribution detection models.
-III-B Datasets for OOD Evaluation
-III-B1 The limits of widely used datasets
-Currently, when addressing the out-of-distribution (OOD) detection problem in 3D object detection, the limitations of existing datasets such as KITTI [Geiger2012CVPR], ONCE [once_ds], and Waymo [Sun_2020_CVPR] become particularly pronounced.
+III A new method for 3D OOD Detection
+While traditional open-set detectors add an extra “unknown” class, we instead view the challenge as an OOD prediction problem.
+Closed-set detectors are optimized for a fixed set of “core” classes, but an OOD framework allows the model to flag deviations from the learned distribution without sacrificing known-class performance.
+In this section, we compare methods that compute “unknown” object scores directly from the detector’s outputs (single-stage methods) with those that add an extra trainable module (two-stage methods).
+We also introduce a novel input representation for two-stage methods.
+III-A Datasets for OOD Evaluation
+III-A1 The limits of widely used datasets
+Currently, when addressing the out-of-distribution (OOD) detection problem in 3D object detection, the limitations of existing datasets such as KITTI [17], ONCE [18], and Waymo [19] become particularly pronounced.
 Indeed, these datasets contain a small number of classes, respectively 5, 5, and 3.
-Methods such as [openset3Dobj], which evaluate on the Van and Truck subset of KITTI, tackle a specific part of the issue, these classes being both vehicles and lacking the high diversity which can arise in real-world scenarios.
-nuScenes [caesar2020nuscenes], while one of the most comprehensive datasets, presents specific challenges critical in the context of OOD detection.
+Methods such as [1], which evaluate on the Van and Truck subset of KITTI, tackle a specific part of the issue, these classes being both vehicles and lacking the high diversity which can arise in real-world scenarios.
+nuScenes [20], while one of the most comprehensive datasets, presents specific challenges critical in the context of OOD detection.
 Firstly, the predefined dataset splits and evaluation metrics in nuScenes might not be sufficient for evaluating OOD performance.
-For instance, only one out of four of the dataset’s scans in its evaluation set contains unknown objects as defined by [revisiting], thus limiting the evaluation interpretation.
+For instance, only one out of four of the dataset’s scans in its evaluation set contains unknown objects as defined by [16], thus limiting the evaluation interpretation.
 nuScenes also suffers from its point cloud sparsity due to its 32-beam LiDAR sensor, which can only be solved by aggregating scans and creating a trailing point cloud for moving objects, hindering detection.
 These findings underscore the need to use a more diverse and dense dataset to effectively tackle the OOD problem in 3D object detection.
-III-B2 Argoverse 2
-In order to establish a robust baseline for out-of-distribution (OOD) detection, we turn to the Argoverse 2 dataset [wilson2021argoverse], which offers several distinct advantages over other datasets such as nuScenes.
+III-A2 Argoverse 2
+In order to establish a robust baseline for out-of-distribution (OOD) detection, we turn to the Argoverse 2 dataset [21], which offers several distinct advantages over other datasets such as nuScenes.
 Argoverse 2 is noted for its high class diversity, encompassing a wide range of object types and scenarios, which is crucial for training models that can generalize well to new, unseen objects.
@@
 The dataset contains scans in the training set and in its validation set.
-For Argoverse 2, we choose the rarest 5% of classes as our open classes as in [revisiting], denoted as set .
-This approach ensures that the chosen open classes are genuinely underrepresented in the training data, providing a challenging test for the OOD detection capabilities of our models.
+We underline its appeal compared to nuScenes in Table I, which shows the percentage of ground truth objects used for metric computation as per the matching and evaluation procedure described in IV-A, i.e.
+detection hits.
+The percentage of total ground truth unknown objects that will be considered for the evaluation is twice as low for nuScenes compared to Argoverse 2, showing more limited interpretability of results.
+| Dataset | Hits % (Open) | Hits % (Closed) |
+| Argoverse2 | 24.1 | 50.2 |
+| nuScenes | 12.9 | 45.0 |
+For Argoverse 2, we choose the rarest 5% of classes as our unknown classes, denoted as set .
+This approach ensures that the chosen unknown classes are genuinely underrepresented in the training data, providing a challenging test for the OOD detection capabilities of our models.
 This selection method helps simulate real-world scenarios where we assume that the most common objects will have been labeled beforehand while certain rare objects are infrequently encountered, making it a valuable benchmark for assessing the robustness of 3D object detection systems.
-The open classes , representing instances in the validation set, are:
+The unknown classes , representing instances in the validation set, are:
 | Articulated bus, Stroller, Motorcycle, Mobile pedestrian crossing | |||
@@
 | Stop sign, Sign, Bus, Truck, Bicycle, Bicyclist, Wheeled device, | |||
-III-B3 Metrics and Matching
-In our evaluation, we use metrics commonly employed in OOD detection, including AUROC, FPR-95, AUPR-E (with open classes as positives), and AUPR-S (with known classes as positives), to assess the model’s ability to distinguish between known and unknown objects.
-We also describe our matching algorithm, detailed in Algorithm 1, which is crucial for computing True Positives (TP), False Positives (FP), True Negatives (TN), and False Negatives (FN).
-This algorithm, which includes essential hyperparameters, plays a significant role in determining the final evaluation metrics, mainly by ignoring any predicted box whose center is not within a fixed range of any non-assigned ground truth.
-III-C Single and Two-Stage Methods for OOD
 Methods to compute ”unknown” object scores can be generally separated into two categories: single-stage methods, which directly use the outputs of a detector model and apply some thresholding or entropy computation, and two-stage methods, which add additional trainable parameters on top of the base detector.
 This latter approach necessitates a retraining phase.
-We illustrate these two families of methods in Figure 2.
-III-C1 Single-Stage Methods
+III-B Single-Stage Methods
 In our evaluation of OOD detection for 3D object detection, we selected several single-stage methods to ensure a comprehensive assessment, i.e.
 methods that directly use the output logits of the detection model.
-The chosen methods include ODIN [odin], MaxLogit [msp], Energy Thresholding [energy], MC Dropout inspired by [mc_dropout], and the Default Score.
-III-C2 Two-Stage Methods
-Like [revisiting], we integrated an additional Multi-Layer Perceptron (MLP) module after the base detector.
+The chosen methods include ODIN [10], MaxLogit [11], Energy Thresholding [12], MC Dropout inspired by [4], and the Default Score.
+III-C Two-Stage Methods
+We introduce our proposed method High-Dimensional OOD 3D Object Detection (HD-OOD3D), which integrates an additional Multi-Layer Perceptron (MLP) module after the base detector.
 This module is specifically trained to output a binary classification confidence score or an ”unknown” confidence score: 0 for known objects and 1 for unknown objects.
-This two-stage solution employs the MLP to process initial detections, decomposed into embeddings, box predictions, and prediction logits and outputs an ”out-of-distribution” score .
+We illustrate the training pipeline of HD-OOD3D in Figure 1.
+This two-stage solution employs the MLP to process initial detections, decomposed into embeddings extracted at the feature map locations of predicted bounding boxes, box predictions , and prediction logits and outputs an ”out-of-distribution” score .
 During inference, if is larger than a given threshold , the object is considered an OOD object.
-III-D Analysis of the Evaluation Protocol for 3D OOD
-We use the ground truth data to train the MLP with the actual annotated ’unknown’ classes, which we will refer to as the ’Oracle MLP.’ We demonstrate that many external factors could impact the performance and resulting metrics, which are not disclosed in [revisiting].
-We identify several key parameters that significantly influence the effectiveness of the Oracle MLP when applied to GT data, allowing us to examine the influence of these parameters independently of the actual performance of the MLP.
-In the first stage, we train a CenterPoint [Yin_2021_CVPR] object detector on the Argoverse 2 dataset [wilson2021argoverse] with classical training regimes [revisiting, Yin_2021_CVPR].
+Unlike MMOOD3D[16], which uses lower-resolution input representation and generates unknowns using a resizing method, we use different input representations and unknown generation methods, the benefits of which we demonstrate in Section III-E and Section V-B respectively.
+In the first stage, we train a CenterPoint [7] object detector on the Argoverse 2 dataset [21] with classical training regimes [16, 7].
 This model obtained a mAP score of 35.2 on the known classes.
-For the two-stage family of methods, we train an MLP after freezing the initial detector Centerpoint.
-We use a 3-layer MLP where we divide the dimensionality by two between every layer, followed by a sigmoid activation function, on top of our frozen pretrained CenterPoint model.
+We freeze this first stage model then add a 3-layer MLP where we divide the dimensionality by two between every layer, followed by a sigmoid activation function.
 We train it for 5 epochs, with a polynomial learning rate using a power of , with an initial learning rate of .
-For ease of reading, we include relevant experiments from the best single-stage model ODIN [odin] (see Section IV-A) and the two-stage Oracle MLP in the following.
-By systematically analyzing the evaluation hyperparameters, we aim to provide a more comprehensive understanding of how different evaluation factors influence the final performance of unknown OOD detection.
-These hyperparameters include:
-III-D1 Evaluation set
-Using the ”Open” Scans subset, which includes only scans with unknown objects, significantly improved the AUPR-E metric from 12.6 to 20.0 for the 2-stage Oracle MLP.
-Due to the known/unknown class imbalance, using all scans adds many more potential false positives without any unknown ground truth, hindering all methods across the board.
-Consequently, our final evaluation used the ”open” subset—only the evaluation scans containing at least one unknown object—to better assess performance.
-III-D2 Distance and Threshold Settings for OOD Detection
-The choices of distance metric and detection module threshold play a critical role in OOD Detection.
-Adjusting the distance threshold for matching predicted boxes to ground truth objects influences the inclusion of predictions in metric calculations.
-Specifically, when using a larger (e.g., 2.0m instead of 0.5m), more predictions, mainly those correctly identifying unknown classes, are included.
-This increase in detection hits—predicted boxes within the range of ground truth—allows for a more comprehensive evaluation of the model’s ability to detect unknown objects.
-However, this also introduces a risk of misassigning predictions to known objects, potentially lowering AUROC and FPR-95 while improving AUPR-E, as this metric directly reflects the precision in detecting unknowns.
-A lower threshold in the detection module (e.g., 0.3) allows more predicted objects to pass through, reducing misses, especially for known objects.
-However, this comes at the cost of an increased likelihood of false positives among the unknown detections.
-A trade-off exists: while a lower threshold increases the detection rate of unknown objects, it also results in more potential false positives.
-Therefore, a threshold of 0.3, paired with a 2.0m distance for box matching, strikes a balance between minimizing misses for known objects and enhancing the detection of unknown objects, as evidenced by the metrics in Table I.
-| Hits (O)% | Hits (C)% | Method | AUROC | FPR-95 | AUPR-E | AUPR-S | ||
-| 0.05 | 0.5m | 22.8 | 49.3 | ODIN | 88.3 | 62.4 | 10.1 | 97.3 |
-| Oracle MLP | 91.0 | 43.7 | 22.2 | 99.8 | ||||
-| 2.0m | 38.9 | 55.4 | ODIN | 82.1 | 78.6 | 9.1 | 97.4 | |
-| Oracle MLP | 84.9 | 62.8 | 17.5 | 99.5 | ||||
-| 0.3 | 0.5m | 14.4 | 45.3 | ODIN | 89.9 | 47.8 | 9.8 | 96.9 |
-| Oracle MLP | 92.3 | 37.4 | 23.2 | 99.9 | ||||
-| 2.0m | 26.2 | 50.6 | ODIN | 85.9 | 61.6 | 9.8 | 97.1 | |
-| Oracle MLP | 88.6 | 51.1 | 20.0 | 99.7 | ||||
-| 0.5 | 0.5m | 7.6 | 39.0 | ODIN | 91.4 | 38.0 | 17.0 | 96.4 |
-| Oracle MLP | 93.6 | 32.6 | 24.8 | 99.9 | ||||
-| 2.0m | 11.7 | 42.4 | ODIN | 87.4 | 48.9 | 14.2 | 96.5 | |
-| Oracle MLP | 91.4 | 38.7 | 22.1 | 99.9 |
-III-D3 Sorting method
-In our experiments, we observed a 2.0 increase in the AUROC metric when sorting predictions by the base detector’s confidence score rather than by the ”unknown” score confidence.
-This suggests that using the original detector’s score is more reliable for prioritizing predictions, especially given the high prediction scores for unknown objects and the non-overlap hypothesis in 3D space.
-From Table I, we notice widely different results from the traditional OOD metrics for any fixed method, depending on the evaluation hyperparameters.
-Notably, they modify the performance differences between methods.
-For our following evaluations, we decided to strike a balance between a high recall of unknowns, which includes noisy predictions and complex examples, and a lower recall, which only isolates well-defined examples.
-We suggest that future works take into account the recall of the base detector and specify these important evaluation hyperparameters.
-IV Comparison of single, two-stage methods and data generation for OOD training
-IV-A Limits of single-stage methods
-We compared multiple single-stage methods for ODD detection on the Argoverse 2 dataset with our evaluation protocol of Section III.
+First, using a methodology we will refer to as ”supervised,” we will leverage the labels of the unknown objects present in the Argoverse 2 dataset.
+This supervised approach allows for a more rigorous comparison of the different methods: one-stage, two-stage, and our HD-OOD3D method against the state of the art.
+In the second part, Section V, we will explore the ”unsupervised” methodology, where we assume we do not have access to the labels of the unknown objects in the Argoverse 2 dataset.
+This will enable us to compare various pseudo-label generation methods for unknown objects.
+III-D Results on unknown learning with single and two-stage methods
+We compared multiple single-stage methods for ODD detection on the Argoverse 2 dataset with our evaluation protocol of Section IV.
+We assess the OOD detector using several metrics: AUROC, which quantifies the overall likelihood that a positive case is ranked above a negative case; FPR-95, which indicates the false positive rate when the true positive rate is set at 95%, balancing sensitivity and specificity; and the AUPR metrics, AUPR-E and AUPR-S, which focus on the model’s precision-recall balance for unknown and known objects respectively.
+Importantly, AUPR-S represents the model performance on known objects.
+If a noticeable downgrade in this metric is noticed, such as under 99, then we consider the base detector’s performance on known objects to be hindered by the additional unknown object classifier module.
 Table II shows the results.
-All single-stage models fail to reach the performance of the 2-stage oracle MLP.
-While the ODIN method achieves the highest AUROC (85.9 on Argoverse 2), it still exhibits a high False Positive Rate at 95% recall (61.6%).
-Energy-based methods, Monte-Carlo Dropout, MaxLogit, and Default Score thresholding yield relatively weaker results.
-These consistently mediocre performances across all methods indicate that single-stage approaches are lacking compared to a model trained directly on unknown object representations.
-Therefore, we experimented with two-stage methods to try and match this oracle.
-| Method | AUROC | FPR-95 | AUPR-E | AUPR-S |
-| Default Score | 76.6 | 86.5 | 4.3 | 99.4 |
-| Energy [energy] | 65.8 | 94.6 | 2.7 | 99.0 |
-| Monte Carlo Dropout [mc_dropout] | 75.4 | 69.7 | 7.0 | 98.9 |
-| MaxLogit [msp] | 76.6 | 86.5 | 4.3 | 98.6 |
-| ODIN [odin] | 85.9 | 61.6 | 9.8 | 97.1 |
-| 2-Stage Oracle MLP | 86.8 | 51.1 | 20.0 | 99.7 |
-IV-B Two-Stage Methods and Input Choices
-The two-stage methods use an MLP for OOD prediction. We first compare different inputs of the MLP using GT data (named ’Oracle MLP’ as in Section III-D):
-high-dimensional feature map (High-Dim): we can use higher-dimension (512) feature map representation from the base detector instead of the final lower-dimension (192) feature maps
+All single-stage models fail to reach the performance of two-stage methods.
+While ODIN achieves the highest AUROC among single-stage approaches (85.9 on Argoverse 2), it still exhibits a high False Positive Rate at 95% recall (61.6%).
+Energy-based methods, Monte-Carlo Dropout, MaxLogit, and Default Score thresholding yield relatively weaker results, with AUROC values ranging from 65.8 to 76.6.
+In contrast, two-stage methods significantly outperform single-stage approaches.
+MMOOD3D reaches an AUROC of 88.6 with a lower FPR-95 of 43.2.
+Our proposed method, HD-OOD3D, achieves competitive results with an AUROC of 86.8 while maintaining a lower FPR-95 (51.1) compared to ODIN.
+Crucially, HD-OOD3D preserves the performance on known objects (AUPR-S = 99.7), whereas MMOOD3D suffers a substantial drop in AUPR-S to 70.5.
+This suggests that HD-OOD3D provides a better balance between detecting unknown objects and preserving the accuracy of the base detector.
+| Type | Method | AUROC | FPR-95 | AUPR-E | AUPR-S |
+| Default Score | 76.6 | 86.5 | 4.3 | 99.4 | |
+| Single | Energy [12] | 65.8 | 94.6 | 2.7 | 99.0 |
+| Stage | Monte Carlo Dropout [4] | 75.4 | 69.7 | 7.0 | 98.9 |
+| MaxLogit [11] | 76.6 | 86.5 | 4.3 | 98.6 | |
+| ODIN [10] | 85.9 | 61.6 | 9.8 | 97.1 | |
+| Two | MMOOD3D [16] | 88.6 | 43.2 | 27.1 | 70.5 |
+| Stage | HD-OOD3D (Ours) | 86.8 | 51.1 | 20.0 | 99.7 |
+III-E Ablation study on HD-OOD3D
+The two-stage methods use an MLP for OOD prediction. We first compare different inputs of the MLP using the labels of unknown objets as previously:
+high-dimensional feature map (High-Dim): we use higher-dimension (512) feature map representation to extract the embeddings from the base detector instead of the final lower-dimension (192) feature maps
 feature interpolation (Interp.): Instead of extracting the feature representation of the base detector’s boxes at the discrete grid value of the feature map, we instead interpolate these features based on the predicted center
@@
 We show our experiments using the ground truth unknown objects for training in Table III.
-The results show that adding a 3x3 pooling layer and a focal loss was beneficial, compared to the work [revisiting], which only uses a binary cross-entropy loss, no pooling, high-dimensional feature maps, and feature interpolation.
-Despite a higher FPR-95, this is the only method that combines a higher precision on known objects and near-perfect precision on closed objects, which was outlined as a primary objective when seeing Open-Set as an OOD problem.
-We highlight this in the table, where using the input processing of [revisiting] leads to a known object precision of 70.5, and ours a 99.7 precision while maintaining an acceptable AUPR-E.
-The pooling smooths out the center interpolation errors from the detection model.
+The results show that adding a 3x3 pooling layer and a focal loss was beneficial, compared to the work [16], which only uses a binary cross-entropy loss, no pooling, high-dimensional feature maps, and feature interpolation.
+Despite a higher FPR-95, this is the only method that combines a higher precision on known objects and near-perfect precision on closed objects.
 For instance, in the Argoverse 2 dataset, a maxpooling filter enabled a feature aggregation operation with a radius of 2.1 meters.
 It substantially improved OOD performance by ensuring a more robust aggregation of surrounding point cloud data.
-While we trained our MLP models using the ground truth unknown objects in the previous section, we cannot access this information when training in real-world scenarios.
-Therefore, we need to generate pseudo-labels to train our model.
-IV-C Methods for Generating Additional Unknown data for training the MLP
-We then trained our MLP to detect OOD objects, using different generation methods of unknown objects, which we use as our ground truth.
-IV-C1 Different Representations of Out-of-Distribution Objects
-We compare different representations of out-of-distribution objects for training, which we create through several means:
-Gaussian noise [du2022vos]: we perturb the extracted embedding of ID objects with Gaussian noise, adding a normal distribution to the extracted embeddings.
-External data: we incorporate ModelNet [modelnet], allowing us to introduce a diverse set of synthetic objects into our training regime, which we consider OOD.
+IV Analysis of the Evaluation Protocol for 3D OOD
+Here, we examine how external factors and hyperparameter choices influence OOD detection performance because these elements directly affect both the quality and reliability of the detection metrics.
+For instance, the characteristics of the dataset—such as point cloud density, class diversity, and the presence of noise—can alter how many true and false positives are identified.
+Similarly, hyperparameters like the distance threshold for matching predictions, the detection score cutoff, and the criteria used for sorting predictions determine which detections are considered valid.
+Even small adjustments in these parameters can lead to significant shifts in performance metrics, affecting the reproducibility and fairness of the evaluation.
+Therefore, we examine and compare these choices.
+IV-A Metrics and Matching
+As previously described, we use the AUROC, FPR-95, AUPR-E (with unknown classes as positives) and AUPR-S (with known classes as positives) to assess the model’s ability to distinguish between known and unknown objects.
+We also describe our matching algorithm, detailed in Algorithm 1, which is crucial for computing True Positives (TP), False Positives (FP), True Negatives (TN), and False Negatives (FN).
+This algorithm, which includes essential hyperparameters, plays a significant role in determining the final evaluation metrics, mainly by ignoring any predicted box whose center is not within a fixed range of any non-assigned ground truth.
+We use the ground truth data to train the MLP with the actual annotated ’unknown’ classes, like in Section III.
+We demonstrate that many external factors could impact the performance and resulting metrics, which are not disclosed in [16].
+We identify several key parameters that significantly influence the effectiveness of HD-OOD3D when applied to GT data, allowing us to examine the influence of these parameters independently of the actual performance of the MLP.
+For the two-stage family of methods, we train an MLP after freezing the initial detector Centerpoint, like in the previous section, and use the same MLP architecture.
+For ease of reading, we include the relevant experiments from our two-stage HD-OOD3D in the following.
+IV-B Hyperparameter Analysis
+By systematically analyzing the evaluation hyperparameters, we aim to provide a more comprehensive understanding of how different evaluation factors influence the final performance of unknown OOD detection.
+These hyperparameters include:
+IV-B1 Evaluation set
+Using the ”Open” Scans subset, which includes only scans with unknown objects, significantly improved the AUPR-E metric from 12.6 to 20.0 for HD-OOD3D.
+Due to the known/unknown class imbalance, using all scans adds many more potential false positives without any unknown ground truth, hindering all methods across the board.
+Consequently, our final evaluation used the ”open” subset—only the evaluation scans containing at least one unknown object—to better assess performance.
+It is important to note that, as shown in Table IV, the chosen hyperparameters do not maximize the performance of HD-OOD3D in isolation, but rather provide a balanced and fair benchmark for all methods.
+IV-B2 Distance and Threshold Settings for OOD Detection
+The choices of distance metric and detection module threshold play a critical role in OOD Detection.
+Adjusting the distance threshold for matching predicted boxes to ground truth objects influences the inclusion of predictions in metric calculations.
+Specifically, when using a larger (e.g., 2.0m instead of 0.5m), more predictions, mainly those correctly identifying unknown classes, are included.
+This increase in detection hits—predicted boxes within the range of ground truth—allows for a more comprehensive evaluation of the model’s ability to detect unknown objects.
+However, this also introduces a risk of misassigning predictions to known objects, potentially lowering AUROC and FPR-95 while improving AUPR-E, as this metric directly reflects the precision in detecting unknowns.
+A lower threshold in the detection module (e.g., 0.3) allows more predicted objects to pass through, reducing misses, especially for known objects.
+However, this comes at the cost of an increased likelihood of false positives among the unknown detections.
+A trade-off exists: while a lower threshold increases the detection rate of unknown objects, it also results in more potential false positives.
+Therefore, a threshold of 0.3, paired with a 2.0m distance for box matching, strikes a balance between minimizing misses for known objects and enhancing the detection of unknown objects, as evidenced by the metrics in Table IV.
+| Hits (U)% | Hits (C)% | AUROC | FPR-95 | AUPR-E | AUPR-S | ||
+| 0.05 | 0.5m | 22.8 | 49.3 | 91.0 | 43.7 | 22.2 | 99.8 |
+| 2.0m | 38.9 | 55.4 | 84.9 | 62.8 | 17.5 | 99.5 | |
+| 0.3 | 0.5m | 14.4 | 45.3 | 92.3 | 37.4 | 23.2 | 99.9 |
+| 2.0m | 26.2 | 50.6 | 88.6 | 51.1 | 20.0 | 99.7 | |
+| 0.5 | 0.5m | 7.6 | 39.0 | 93.6 | 32.6 | 24.8 | 99.9 |
+| 2.0m | 11.7 | 42.4 | 91.4 | 38.7 | 22.1 | 99.9 |
+IV-B3 Sorting method
+In our experiments, we observed a 2.0 increase in the AUROC metric when sorting predictions by the base detector’s confidence score rather than by the ”unknown” score confidence.
+This suggests that using the original detector’s score is more reliable for prioritizing predictions, especially given the high prediction scores for unknown objects and the non-overlap hypothesis in 3D space.
+From Table IV, we notice widely different results from the traditional OOD metrics for any fixed method, depending on the evaluation hyperparameters.
+The choice of detection score threshold and distance threshold is crucial for balancing the precision and recall of OOD detection.
+Both and determine how detections are matched to ther ground truth: larger threshold includes more detections but risk misassigning predictions.
+Rather than focusing on optimizing the performance of our method, our primary goal was to refine the evaluation protocol to ensure a fair and robust comparison.
+Through systematic evaluation, we found that setting and m provides the best trade-off, maximizing recall for OOD objects while minimizing degradation on known classes.
+These choices allow for a more reliable assessment of OOD detection performance across different methods.
+We suggest that future works consider the recall of the base detector and explicitly specify these crucial evaluation hyperparameters to facilitate reproducibility and fair benchmarking.
+V Methods for Generating Additional Unknown data for training the MLP
+In the previous sections, our MLP models were trained using ground truth labels for unknown objects.
+However, in real-world scenarios, manually annotating every unknown object is infeasible due to the diversity and unpredictability of OOD instances.
+Unlike in controlled datasets like Argoverse 2, where unknown objects can be labeled explicitly, a practical deployment scenario requires detecting OOD objects without prior supervision.
+To address this challenge, we propose generating pseudo-labels for unknown objects, enabling training without explicit ground truth annotations.
+V-A Different Representations of Out-of-Distribution Objects
+In contrast to the single-generation approach used in MMOOD3D [16], we propose several alternative methods to synthesize out-of-distribution (OOD) objects, specifically a Top-K labelling approach.
+We compare these different representations of out-of-distribution objects for training, which we create through several means:
+Gaussian noise [9]: we perturb the extracted embedding of ID objects with Gaussian noise, adding a normal distribution to the extracted embeddings.
+External data: we incorporate ModelNet [22], allowing us to introduce a diverse set of synthetic objects into our training regime, which we consider OOD.
 We insert between 15 and 25 objects, sampled randomly from the {’bench’, ’bookshelf’, ’chair’, ’cone’, ’desk’, ’flower pot’, ’lamp’, ’mantel’, ’night stand’, ’plant’, ’vase’, ’wardrobe’} class subset of ModelNet, and normalize their size to 1.
 We then sample 200 points at the surface of each object, followed by a grid sampling where we keep between 1 and 3 points with a grid size which we also randomize between 5 and 10, apply an intensity value to the points based on the intensity statistics of Argoverse 2, and finally resize the objects with a global scaling factor between 1 and 4.
-PointMixup [pointmixup]: we use this method to mix point clouds of different objects through point matching and interpolation to create new, hybrid representations.
+PointMixup [23]: we use this method to mix point clouds of different objects through point matching and interpolation to create new, hybrid representations.
 We expect objects created in such a fashion to contain local features present in both ID and OOD objects while creating novel appearances.
-Since [pointmixup] was designed for data augmentation, we chose different interpolation hyperparameters after testing variations.
+Since [23] was designed for data augmentation, we chose different interpolation hyperparameters after testing variations.
 Each object with at least 5 points has a 20% probability of being replaced by a new version of that object generated as an interpolation between it and another object with at least 5 points, with interpolation factors randomly chosen in the range.
-Resizing: create out-of-distribution objects by resizing in-distribution (ID) known objects to modify their scales along all three axes, similarly to [revisiting].
-Top-K autolabelling approach [towardsopen]: the top-K most confident predictions from the base detector, which don’t overlap with ground truth ID objects, are used as OOD objects.
+Resizing: create out-of-distribution objects by resizing in-distribution (ID) known objects to modify their scales along all three axes, similarly to [16].
+Each known object with at least 5 points has a 50% probability of being resized, where for each axis, we randomly apply a scaling factor between and .
+Our 3D Top-K autolabelling : inpired by [3] for open 2d objet detection in images, we propose in 3D to use the top-K most confident predictions from the base detector, which don’t overlap with ground truth known objects, as OOD objects.
 In this approach, we keep the scans from the datasets containing real OOD distribution, as we expect these to become autolabelled.
-Since other methods do not rely on real OOD data for training the MLP, we remove those scans for other methods.
+Since other methods do not rely on real OOD data for training the MLP, we remove those LiDAR scans for other methods.
 We illustrate these different generation methods in Figure 2 and compare these various representations of OOD objects and their impact on the training and performance of our additional module on real validation OOD data.
-IV-C2 Results
-The results of our evaluation on the Argoverse 2 dataset, presented in Table IV, highlight the discrepancies in the MLP performance based on different generation methods of unknowns.
+V-B Results
+The results of our evaluation on the Argoverse 2 dataset, presented in Table V, highlight the discrepancies in the MLP performance based on different generation methods of unknowns.
 The Resizing method obtained an AUROC of 77.2, with a relatively low AUPR-E of 6.3.
@@
 The Gaussian Noise method performed poorly across all metrics, with an AUROC of 66.5 and a very high FPR-95 of 84.5, making it only slightly better than random chance.
-| Method | AUROC | FPR-95 | AUPR-E | AUPR-S |
-| Gaussian Noise [du2022vos] | 66.5 | 84.5 | 3.7 | 98.9 |
+| Unsupervised Unknowns | AUROC | FPR-95 | AUPR-E | AUPR-S |
+| Gaussian Noise [9] | 66.5 | 84.5 | 3.7 | 98.9 |
 | ModelNet Injection† | 70.6 | 78.7 | 5.0 | 71.1 |
-| PointMixUp†[pointmixup] | 72.1 | 72.4 | 7.4 | 99.1 |
-| Resizing [revisiting] | 77.2 | 75.3 | 6.3 | 99.4 |
-| Top-K Labelling (K=5) [towardsopen] | 81.9 | 68.2 | 8.8 | 99.6 |
-| 2-Stage Oracle MLP | 86.8 | 51.1 | 20.0 | 99.7 |
+| PointMixUp†[23] | 72.1 | 72.4 | 7.4 | 99.1 |
+| Resizing from MMOOD3D[16] | 77.2 | 75.3 | 6.3 | 99.4 |
+| Top-5 Labelling (Ours) | 81.9 | 68.2 | 8.8 | 99.6 |
+| GT Unknowns | 86.8 | 51.1 | 20.0 | 99.7 |
 ModelNet Injection, which incorporates synthetic objects from the ModelNet dataset as OOD data, showed better performance with an AUROC of 70.6 and an AUPR-E of 5.0.
-We believe the difficulties in OOD detection for this method stem from the synthetic nature of the dataset, which suggests that while other works suggest using insertions of instances as unknowns [UFO], one has to be careful about the quality of these instances and the domain differences between these and the real unknowns.
+We believe the difficulties in OOD detection for this method stem from the synthetic nature of the dataset, which suggests that while this method might be interesting [14], one has to be careful about the quality of these instances and the domain differences between these and the real unknowns.
 PointMixUp struggled with a high FPR-95 of 77.6 and a low AUPR-E of 4.6, indicating challenges in detecting unknown objects despite the novel hybrid representations.
-The Top-K Labelling (K=5) method, which labels the top K most confident predictions not overlapping with ground truth ID objects as OOD, emerged as a notable performer.
-It achieved an AUROC of 81.9 and the highest AUPR-S of 99.6, demonstrating exceptional identification of known objects.
+The Top-K Labelling (K=5) method, which labels the top 5 most confident predictions not overlapping with ground truth known objects as OOD, emerged as a notable performer.
+It achieved an AUROC of 81.9 and the highest AUPR-S of 99.6, demonstrating good identification of known objects.
 Despite its higher FPR-95 of 68.2, Top-K Labelling showed promise in balancing the detection of known and unknown objects by leveraging real-world OOD distributions in the dataset.
-We experimented with different K values (K= ) and found that K=5, close to the average number of open class instances per frame, performed best across all metrics.
-Overall, the remaining gap between the two-stage oracle MLP and this highest-performing unknown generation process suggests that future works should focus on improving this pseudo-labelling process.
-V CONCLUSIONS
-Our work redefines the open-set 3D Object Detection problem in LiDAR data as an Out-Of-Distribution problem.
-Through this lens, OOD 3D object detection adds relevant information with regard to traditional object detection.
-In the process, we established a rigorous evaluation protocol and identified key evaluation parameters that we suggest highlighting in future works.
-Additionally, we show that two-stage methods show promise for OOD 3D Object Detection and underline the potential of autolabelling methods.
+We experimented with different K values (K= ) and found that K=5, close to the average number of unknown class instances per frame, performed best across all metrics.
+Overall, the remaining gap between the two-stage oracle MLP and this highest-performing unknown generation method suggests that future works should focus on improving this pseudo-labelling process.
+VI CONCLUSIONS
+In this work, we presented a new two-stage method for detecting unknown objects called HD-OOD3D.
+By exploiting a sufficiently large dataset such as Argoverse 2 to train the unknowns in a supervised manner, we demonstrated the superiority of two-stage methods over single-stage methods, particularly our proposed HD-OOD3D, which utilizes more input representations for better object categorization.
+We then broke down the classic evaluation protocol for unknown objects to highlight the importance of certain hyperparameter choices for a fair assessment of methods.
+Finally, scaling the learning of unknown objects can only be achieved through unsupervised training.
+We compared different approaches for generating unknowns by using only the annotations of known objects from Argoverse 2.
+We showed that top-5 auto-labelling offers more promising performance perspectives compared to resizing.
+Ultimately, the task of detecting unknown objects remains a challenging one, with current performance far from what would enable deployment on robotics platforms, making it a field that still requires further research to reach maturity.